# Stress-Testing Long-Context Language Models with Lifelong ICL and Task Haystack

### v1

@@ Abstract
We introduce Lifelong ICL, a problem setting that challenges long-context language models (LMs) to learn from a sequence of language tasks through in-context learning (ICL).Equal Contribution. We further introduce Task Haystack, an evaluation suite dedicated to assessing and diagnosing how long-context LMs utilizes contexts in Lifelong ICL. When given a task instruction and test inputs, long-context LMs are expected to leverage the relevant demonstrations in the Lifelong ICL prompt, avoid distraction and interference from other tasks, and achieve test accuracies that are not significantly worse than the Single-task ICL baseline.
Task Haystack draws inspiration from the widely-adopted “needle-in-a-haystack” (NIAH) evaluation, but presents new and unique challenges. It demands that models (1) utilize the contexts with deeper understanding, rather than resorting to simple copying and pasting; (2) navigate through long streams of evolving topics and tasks, which closely approximates the complexities of real-world usage of long-context LMs. Additionally, Task Haystack inherits the controllability aspect of NIAH, providing model developers with tools and visualizations to identify model vulnerabilities effectively.
@@ 1 Introduction
Recent advances in model architecture (Han et al., 2024; Su et al., 2024), hardware-aware optimization (Dao et al., 2022; Liu et al., 2023), training procedure (Tworkowski et al., 2023; Liu et al., 2024a), and data engineering (Fu et al., 2024; An et al., 2024) have enabled large language models (LLMs) to handle extended contexts, reaching up to 32k tokens or even millions (Gemini Team Google, 2024; Anthropic, 2024). These advancements have opened up new opportunities and potential use cases for LLMs. However, while long-context LM development strides forward, suitable evaluation methods have not kept pace. Systematically evaluating long-context LMs’ ability to leverage such long contexts remains an open challenge.
Current evaluation approaches fall into two major categories. The first involves constructing benchmarks with real-world long-context tasks (Shaham et al., 2022, 2023). While valuable, creating these benchmarks is time-consuming and particularly challenging when scaling the input context length to millions of tokens. The second approach employs synthetic evaluations like the "needle-in-a-haystack" (NIAH) test (Kamradt, 2023) or key-value retrieval tests (Liu et al., 2024b). For example, in the NIAH evaluation, a piece of information (“The special magic number is 12345”) is planted in a haystack of irrelevant contexts (Paul Graham essays; Graham 2024) and the model is evaluated on answering a question about the information (“What’s the special magic number?”). Although useful for initial assessment, these tests primarily measure simple copying-and-pasting capabilities and fail to capture how models perform when deeper context understanding is crucial.
@@
Task Haystack presents unique challenges not fully covered by existing benchmarks. Firstly, Task Haystack demands a deeper comprehension of the relevant context for accurate predictions. This goes beyond simple retrieval capabilities tested by NIAH-style benchmarks, which often rely on basic copying and pasting. Secondly, Task Haystack features high information density, meaning that every piece of information in the context might be crucial for successful prediction at test time. This differs from evaluation suites in which the important information (“needle”) is positioned conspicuously, allowing models to exploit shortcuts (Anthropic, 2024). Thirdly, the lifelong task stream closely mirrors real-world applications, such as a 24/7 personal assistant, where models encounter shifting topics (Zhao et al., 2024) and may need to resume from earlier threads in the context.
We extensively evaluate 12 long-context models on Task Haystack. While all models achieve near-perfect scores on the original NIAH test, none reach satisfactory performance on our proposed evaluation. Among the compared models, GPT-4o emerged as the top performer, averaging an 85% pass rate and surpassing open models by a large margin. To understand the root causes behind these failure cases, we conduct controlled experiments that isolate factors like recency bias (models favoring information at the end of the context) and distractability (models getting distracted by irrelevant information). The results confirm that both factors contribute to performance degradation on Task Haystack. Additionally, model performance drops when instructions are paraphrased at test time and when few-shot ICL demonstrations of a single task are repeated multiple times. These observations highlight the limitations of current long-context LMs in terms of robustness, instruction understanding, and context utilization.
@@ 2 Related Work
Long-Context LM Benchmarks.
Existing benchmarks for evaluating long-context models can be categorized into realistic and synthetic ones. Realistic benchmarks, exemplified by (Zero)SCROLLS (Shaham et al., 2022, 2023), comprise tasks that require processing long inputs with long-range dependencies. These tasks are typically sourced from established datasets and include various task types such as summarization and question answering. In the category of synthetic benchmarks, the needle-in-a-haystack (NIAH) (Kamradt, 2023) evaluation is widely adopted for evaluating context utilization (Gemini Team Google, 2024; Anthropic, 2024; Liu et al., 2024a; Fu et al., 2024). Ruler (Hsieh et al., 2024) expands on the NIAH test with multi-key and multi-value retrieval, and adds two new tasks that involve multi-hop tracing and aggregation. Hybrid benchmarks is a middle-field that incorporate both realistic and synthetic elements. An example is LongBench (Bai et al., 2023), which includes synthetic tasks based on realistic text, such as counting unique passages appearing in the context. Our proposed Task Haystack can be considered as a hybrid benchmark, with a realistic touch as it closely mirrors the real-world challenge of navigating through evolving topics and tasks.
@@
Several recent works have explored in-context learning with long-context LMs by scaling the number of training examples (i.e., shots). Bertsch et al. (2024) conducted a systematic study of long-context ICL with up to 2000 shots, demonstrating many-shot ICL as a competitive alternative to retrieval-based ICL and fine-tuning. Additionally, it offers the advantage of caching demonstrations at inference time, unlike instance-level retrieval methods. While Bertsch et al. (2024) focus on classification tasks, Agarwal et al. (2024) showed the effectiveness of many-shot ICL on generative and reasoning tasks, and established new state-of-the-art results on practical applications such as low-resource translation with the Gemini 1.5 Pro model. However, there are still limitations to many-shot ICL. Li et al. (2024) introduce LongICLBench, a suite of 6 classification tasks with many (20+) classes, and find that current long-context LMs still struggle with these tasks. Orthogonal to this line of work on scaling number of examples for one single task, we focus on scaling the number of tasks in our Lifelong ICL setting.
@@ 3 Problem Setting
In the following, we establish the notations and the problem setting of Lifelong ICL in §3.1. We will begin by defining notations of in-context learning (ICL) of one single task $T$. We will then build upon these foundations and introduce lifelong ICL with a collection of tasks $\mathcal{T}$. In §3.2, we further introduce our Task Haystack evaluation protocol, provide the definition of pass rate - the key evaluation metric, and describe our strategies to account for the instabilities in ICL experiments.
@@ 3.1 Lifelong ICL
The definition above introduces how ICL is performed with one single task $T$. In Lifelong ICL, an LM is expected to learn from a collection of $n$ tasks, denoted as $\mathcal{T}=\{T_{i}\}_{i=1}^{n}$. To enable this, we first create a random permutation $a=(a_{1},a_{2},\dots,a_{n})$, thus the tasks in $\mathcal{T}$ will be ordered as $(T_{a_{1}},T_{a_{2}},\dots,T_{a_{n}})$. For example, when $n=3$, one possible permuatation $a$ is $(3,1,2)$, so that the tasks are ordered as $(T_{3},T_{1},T_{2})$.
@@ 3.2 Task Haystack
For a test task $T_{a_{i}}$, we anticipate that long-context LMs can effectively utilize the in-context examples of that task, i.e., $p_{a_{i}}$, which is a substring of the lifelong ICL prompt $p_{l}\oplus d_{a_{i}}$. To evaluate this, we compare the model performance on task $T_{a_{i}}$ when conditioning on $p_{l}\oplus d_{a_{i}}$ and $p_{a_{i}}$, and expect the former to be not significantly worse than the latter. In other words, the Single-task ICL prompt $p_{a_{i}}$ is the “needle” in the Lifelong ICL prompt $p_{l}$ (i.e., the “task haystack”). [^2] While the absolute performance of Lifelong ICL will be influenced by various factors, such as the LM’s core capabilities, its parametric knowledge, the prompt template, or the selection of ICL examples, it is reasonable to make a comparative assumption that the performance of Lifelong ICL should not be worse than Single-task ICL for the same model. Additionally, since the Single-task ICL prompt $p_{a_{i}}$ is a substring of the Lifelong ICL $p_{l}$, the quality of the ICL examples are controlled to be the same.
@@
For an overall measurement, we introduce an overall pass rate. For each permutation $a$ and each task $T_{a_{i}}$, we will get two groups of performance metrics, when using sing-task ICL and lifelong ICL respectively. Each group contains 5 metrics, corresponding to the 5 randomly-sampled few-shot training set $D_{train}$. The model passes the test (i.e., scores 1) when the the lifelong ICL group is not significantly worse than the single-task ICL group, captured by a two-sided t-test, and the model scores 0 otherwise. The overall pass rate will be computed by averaging over the different permutations and tasks. See Fig. 6 for an illustration.
@@ 4 Experiment Details
While the problem setting in §3 is generic and admits any language task, in this work we instantiate the setting with a narrower task distribution for initial exploration. Our key considerations include [^4] We discuss the limitations of these design choices in §6. We invite future work to improve upon our work and address these limitations.:
  • We focus on classification tasks, as they allow standardized evaluation. Additionally, a large body of past work investigates ICL empirically or mechanistically using classification tasks (Chang and Jia, 2023; Wang et al., 2023; Chang et al., 2024).
@@
We evaluate ten open-weight long-context LMs on Task Haystack: Mistral-7B (32k) (Jiang et al., 2023), FILM-7B (32k) (An et al., 2024), Llama2-7B (32k) (TogetherAI, 2024), Llama2-7B (80k) (Fu et al., 2024), Llama3-8B/70B (1048k) (GradientAI, 2024a, b), Yi-6B/9B/34B (200k) (01.AI et al., 2024), and Command-R-35B (128k) (Cohere for AI, 2024). These models represent various long-context modeling techniques, model size, and base pre-trained models. We provide the detailed descriptions of these models in Table 5. Additionally, we evaluate two closed models, GPT-3.5-Turbo (16k) and GPT-4o (128k) from OpenAI.
@@
We consider creating long contexts with two strategies: (1) Scale-Shot: scaling the number of in-context examples ($n_{shot}$); (2) Scale-Task: scaling the number of tasks ($n_{task}$). In the first setting, we fix $n_{task}=16$ and experiment with $n_{shot}\in\{1,2,3,4,5,6,7,8\}$. We use the 16 tasks listed in Table 1. In the second setting, we fix $n_{shot}=2$ and experiment with $n_{task}\in\{8,16,24,32,40,48,56,64\}$. Note that to ensure the in-context examples are balanced and every class is covered, $n_{shot}=2$ refers to using 2 examples per class for in-context learning. In both scaling settings, we are able to effectively create contexts of sizes ranging from 4k to 32k tokens. [^5] It is possible to further increase the context length, e.g., reaching 128k tokens with 64 tasks and 8 shots. Due to compute constraints, we limit the context length to 32k in this work.
We defer additional implementation and engineering details in Appendix A.4.
@@ 5.1 Main Results
Table 2: Main Results: Fixing 16 Tasks, Scaling the Number of Shots. “s-acc” stands for single-task ICL accuracy averaged over all 16 tasks, and “l-acc” stands for lifelong ICL accuracy. “pass” represents the pass rate defined in §3.2, i.e., percentage of cases that lifelong ICL is not significantly worse than single-task ICL among 5 random samples of few-shot training sets. l-acc is expected to be not worse than s-acc, and the pass rate is expected to be close to 100%.
@@
Llama2-7B (32k) 61.9 69.8 63.3 77.5 72.8 64.5 53.8 75.6 63.0 41.2 78.0 - -
Llama2-7B (80k) 38.4 47.6 60.0 100.0 49.8 60.2 100.0 56.3 62.3 96.3 59.8 61.5 76.3
Llama3-8B (1048k) 51.2 65.5 68.1 78.8 70.0 69.1 76.2 71.5 70.1 71.3 73.6 70.1 57.5
Llama3-70B (1048k) 60.7 79.1 72.9 68.8 79.0 74.4 50.0 80.3 75.3 57.5 81.7 75.7 51.2
@@
We present the aggregated results (mean accuracy and overall pass rate) of the Scale-Shot setting in Table 2 and the results of the Scale-Task setting in Table 6. The overall pass rates fall below 90% in 42 out of 46 cases reported in Table 2 and in 37 out of 40 cases in Table 6. When scaling to 32k context with 8 shots across 16 tasks, 8 out of the 9 open-weight models achieve pass rates lower than 60%, suggesting that these models are still far from fully utilizing and flexibly conditioning on the provided context. In the most extreme case, Yi-6B (200k) achieves a pass rate of merely 38.8% in the 8-shot (32k) setting.
While model developers commonly use near-perfect needle-in-a-haystack results as evidence of successful long context utilization (01.AI et al., 2024; GradientAI, 2024b, a), our Task Haystack exposes previously unknown limitations of these models and suggest that these models are far from perfect when contextual understanding is required.
@@
To have a holistic view on this, we visualize the results from our Scale-Shot experiments by plotting the Lifelong ICL accuracy and pass rate as a function of Single-task ICL accuracy in Fig. 3. For nearly all models, pass rates decrease when the context length increases, highlighting that while these models are able to take in long context as inputs, they are not necessarily utilizing them effectively. For model-wise comparison, GPT-4o outperforms all other models significantly, in terms of both the ICL accuracy and the pass rate. Mistral-7B and FILM-7B (fine-tuned from Mistral-7B) achieves the strongest performance among open-weight models.
One outlier that we notice is the Llama2-7B (80k) model (Fu et al., 2024), which achieves low ICL accuracies but high pass rates. We notice that this model is trained on language modeling objectives without further instruction tuning or RLHF, which may be the reason behind this trend. This observation also suggests that the pass rates should always be considered together with metrics representing the model’s core capabilities.
@@
Task Haystack supports straightforward visualization for diagnosing model vulnerabilities. In Fig. 3 we present the results of Task Haystack (Scale-Shot Setting) in a way similar to the original needle-in-a-haystack (NIAH) eval. While FILM-7B achieves near-perfect results in the original NIAH eval, Fig. 3 suggests that it’s vulnerable when the context length exceeds 12k, particularly when the relevant information appears in the first 75% of the context window. We include NIAH-style visualizations for all compared models in Fig. 11-20. In addition, we provide examples of aggregating results by permutations, by depth in the context, and by task in Fig. 23-27 and further discuss our findings in §D.2.

### v2

@@ Abstract
We introduce Lifelong ICL, a problem setting that challenges long-context language models (LMs) to learn a sequence of language tasks through in-context learning (ICL).Equal Contribution. We further introduce Task Haystack, an evaluation suite dedicated to assessing and diagnosing how long-context LMs utilizes contexts in Lifelong ICL. When given a task instruction and test inputs, long-context LMs are expected to leverage the relevant demonstrations in the Lifelong ICL prompt, avoid distraction and interference from other tasks, and achieve test accuracies that are not significantly worse than those of the Single-task ICL baseline.
Task Haystack draws inspiration from the widely-adopted “needle-in-a-haystack” (NIAH) evaluation, but presents distinct new challenges. It requires models (1) to utilize the contexts at a deeper level, rather than resorting to simple copying and pasting; (2) to navigate through long streams of evolving topics and tasks, proxying the complexities and dynamism of contexts in real-world scenarios. Additionally, Task Haystack inherits the controllability of NIAH, providing model developers with tools and visualizations to identify model vulnerabilities effectively.
@@ 1 Introduction
Recent advances in model architecture (Han et al., 2024; Su et al., 2024), hardware-aware optimization (Dao et al., 2022; Liu et al., 2024b), training procedure (Tworkowski et al., 2023; Liu et al., 2024a), and data engineering (Fu et al., 2024; An et al., 2024) have enabled large language models (LLMs) to handle extended contexts, reaching up to 32 thousand tokens or even millions (Gemini Team, 2024; Anthropic, 2024). These advancements have opened up new opportunities and potential use cases for LLMs. However, while long-context LM development strides forward, effective evaluation methods have not kept pace. Systematically evaluating long-context LMs’ ability to leverage such long contexts remains an open challenge.
Current evaluation approaches fall into two major categories. The first involves constructing benchmarks with real-world long-context tasks (Shaham et al., 2022, 2023). While valuable, creating these benchmarks is time-consuming and particularly challenging when scaling the input context length to millions of tokens. The second approach employs synthetic evaluations like the “needle-in-a-haystack” (NIAH) test (Kamradt, 2023) or key-value retrieval tests (Liu et al., 2024c). For example, in the NIAH evaluation, a piece of information (“The special magic number is 12345”) is planted in a haystack of irrelevant contexts (Paul Graham essays; Graham 2024) and the model is evaluated on answering a question about the information (“What’s the special magic number?”). Although useful for initial assessment, these tests primarily measure simple copying-and-pasting capabilities and fail to capture whether models are able to utilize the context at a deeper level.
@@
Task Haystack presents unique challenges not fully covered by existing benchmarks. Firstly, Task Haystack requires deeper understanding of the relevant context for accurate predictions. This goes beyond simple retrieval capabilities tested by NIAH-style benchmarks, which often rely on basic copying and pasting. Secondly, Task Haystack features high information density, meaning that every piece of information in the context might be crucial for successful prediction at test time. This differs from evaluation suites in which the important information (“needle”) is positioned conspicuously, allowing models to exploit shortcuts (Anthropic, 2024). Thirdly, existing benchmarks fall short in capturing the dynamics of shifting topics within the context (Zhao et al., 2024), which can pose challenges in real-world applications of long-context models—such as a 24/7 personal assistant that must resume previous conversations amid a long, evolving stream of topics. While not fully realistic, Task Haystack serves as a useful proxy for evaluating this aspect.
We extensively evaluate 14 long-context models on Task Haystack. While all models achieve near-perfect scores on the original NIAH test, none reach satisfactory performance on our proposed evaluation. Among the compared models, GPT-4o and Gemini-1.5-Flash lead with an average pass rate of 85%, significantly outperforming most open-weight models. Llama-3.1-70B, the best-performing open-weight model, follows closely with an average pass rate of 80%. To understand the root causes behind these failure cases, we conduct controlled experiments that isolate factors like recency bias (models favoring information at the end of the context) and distractability (models getting distracted by irrelevant information). The results confirm that both factors contribute to performance degradation on Task Haystack. Additionally, we find that model performance declines when instructions are paraphrased at test time and when few-shot ICL demonstrations of a single task are repeated multiple times. These observations highlight the limitations of current long-context LMs in terms of robustness, instruction understanding, and context utilization.
@@ 2 Related Work
Long-Context LM Evaluation.
Early studies on long-context modeling primarily rely on perplexity-based evaluations (Beltagy et al., 2020; Press et al., 2022). Subsequent research has indicated that such evaluation is limited in reflecting a model’s effectiveness in downstream applications (Sun et al., 2021; Hu et al., 2024). Recent efforts have led to the development of comprehensive benchmarks for evaluating long-context models, which can be divided into realistic and synthetic categories. Realistic benchmarks, exemplified by (Zero)SCROLLS (Shaham et al., 2022, 2023), comprise tasks that require processing long inputs collected from real-world scenarios. These tasks are typically sourced from established datasets and include various task types such as summarization and question answering, or developed from inherently lengthy corpus such as novel (Zhang et al., 2024), grammar books (Tanzer et al., 2024) and code repository (Jimenez et al., 2024). In the category of synthetic benchmarks, the needle-in-a-haystack (NIAH) (Kamradt, 2023) evaluation is widely adopted for evaluating context utilization (Gemini Team, 2024; Anthropic, 2024; Liu et al., 2024a; Fu et al., 2024; Levy et al., 2024, i.a.). Ruler (Hsieh et al., 2024) expands on the NIAH test with multi-key and multi-value retrieval, and adds two new tasks that involve multi-hop tracing and aggregation. Hybrid benchmarks is a middle-field that incorporate both realistic and synthetic elements. An example is LongBench (Bai et al., 2024), which includes synthetic tasks based on realistic text, such as counting unique passages appearing in the context. Our proposed Task Haystack can be seen as a hybrid benchmark, with a realistic touch as (1) it is built upon realistic language tasks; (2) it proximates the challenge of navigating through evolving topics and tasks.
@@
Several recent works have explored in-context learning with long-context LMs by scaling the number of training examples (i.e., shots). Bertsch et al. (2024) conducted a systematic study of long-context ICL with up to 2,000 shots, demonstrating many-shot ICL as a competitive alternative to retrieval-based ICL and fine-tuning. Additionally, it offers the advantage of caching demonstrations at inference time, unlike instance-level retrieval methods. While Bertsch et al. (2024) focus on classification tasks, Agarwal et al. (2024) showed the effectiveness of many-shot ICL on generative and reasoning tasks, and established new state-of-the-art results on practical applications such as low-resource translation with the Gemini 1.5 Pro model. However, there are still limitations to many-shot ICL. Li et al. (2024) introduce LongICLBench, a suite of 6 classification tasks with many (20+) classes, and find that current long-context LMs still struggle with these tasks. Orthogonal to this line of work on scaling number of examples for one single task, we focus on scaling the number of tasks in our Lifelong ICL setting.
@@ 3 Problem Setting
In the following, we establish the notations and the problem setting of Lifelong ICL in §3.1. We will begin by defining notations of in-context learning (ICL) of one single task $T$. We will then build upon these foundations and introduce Lifelong ICL with a collection of tasks $\mathcal{T}$. In §3.2, we further introduce our Task Haystack evaluation protocol, provide the definition of the key metric named “pass rate,” and describe our strategies to account for the instabilities in ICL experiments.
@@ 3.1 Lifelong ICL
The definition above introduces how ICL is performed with one single task $T$. In Lifelong ICL, an LM is expected to learn from a collection of $n$ tasks, denoted as $\mathcal{T}=\{T_{i}\}_{i=1}^{n}$. To enable this, we first create a random permutation $a=(a_{1},a_{2},\dots,a_{n})$, thus the tasks in $\mathcal{T}$ will be ordered as $(T_{a_{1}},T_{a_{2}},\dots,T_{a_{n}})$. For example, when $n=3$, one possible permutation $a$ is $(3,1,2)$, so that the tasks are ordered as $(T_{3},T_{1},T_{2})$.
@@ 3.2 Task Haystack
For a test task $T_{a_{i}}$, we anticipate that long-context LMs can effectively utilize the in-context examples of that task, i.e., $p_{a_{i}}$, which is a substring of the Lifelong ICL prompt $p_{l}\oplus d_{a_{i}}$. To evaluate this, we compare the model performance on task $T_{a_{i}}$ when conditioning on $p_{l}\oplus d_{a_{i}}$ and $p_{a_{i}}$, and expect the former to be not significantly worse than the latter. In other words, the Single-task ICL prompt $p_{a_{i}}$ is the “needle” in the Lifelong ICL prompt $p_{l}$ (i.e., the “task haystack”). [^2] While the absolute performance in Lifelong ICL will be influenced by various factors, such as the LM’s core capabilities, its parametric knowledge, the prompt template, or the selection of ICL examples, it is reasonable to make a comparative assumption that the performance of Lifelong ICL should not be worse than Single-task ICL for the same model. Additionally, since the Single-task ICL prompt $p_{a_{i}}$ is a substring of the Lifelong ICL $p_{l}$, the quality of the ICL examples are controlled to be the same.
@@
For an overall measurement, we introduce an overall pass rate. For each permutation $a$ and each task $T_{a_{i}}$, we will get two groups of performance metrics, when using Single-task ICL and Lifelong ICL respectively. Each group contains 5 metrics, corresponding to the 5 randomly-sampled few-shot training set $D_{train}$. The model passes the test (i.e., scores 1) when the the Lifelong ICL group is not significantly worse than the Single-task ICL group, captured by a two-sided t-test with $p=0.05$. The model scores 0 otherwise. The overall pass rate will be computed by averaging the scores over the different permutations and tasks. We provide more details of the definition and discuss its limitations in §A.3.
@@ 4 Experiment Details
While the problem setting in §3 is generic and admits any language task, in this work we instantiate the setting with a narrower task distribution for initial exploration. Our key considerations include: [^4] We discuss the limitations of these design choices in §6. We invite future work to improve upon our work and address these limitations.
  • We focus on classification tasks, as they allow standardized evaluation. Additionally, a large body of past work investigates ICL empirically or mechanistically using classification tasks (Halawi et al., 2023; Chang and Jia, 2023; Wang et al., 2023; Chang et al., 2024, i.a.).
@@
We evaluate eleven open-weight long-context LMs on Task Haystack: Mistral-7B (32k) (Jiang et al., 2023), FILM-7B (32k) (An et al., 2024), Llama-2-7B (32k) (TogetherAI, 2024), Llama-2-7B (80k) (Fu et al., 2024), Llama-3-8B/70B (1048k) (GradientAI, 2024a, b), Llama-3.1-70B (128k) (Dubey et al., 2024), Yi-6B/9B/34B (200k) (01.AI et al., 2024), and Command-R-35B (128k) (Cohere for AI, 2024). These models represent various long-context modeling techniques, model size, and base pre-trained models. Additionally, we evaluate three closed models, GPT-3.5-Turbo (16k) and GPT-4o (128k) from OpenAI, and Gemini-1.5-Flash (1048k) from Google DeepMind (Gemini Team, 2024). We provide the detailed descriptions of these models in Table 5 in §A.1.
@@
We consider creating long contexts with two strategies: (1) Scale-Shot: scaling the number of in-context examples ($n_{shot}$); (2) Scale-Task: scaling the number of tasks ($n_{task}$). In the first setting, we fix $n_{task}=16$ and experiment with $n_{shot}\in\{1,2,3,4,5,6,7,8\}$. We use the 16 tasks listed in Table 1 in the main body of the paper. [^5] Following reviewer feedback, we create a separate subset of 16 tasks with permissive licenses, and report the results on selected models in §B.2. We recommend using this subset for future benchmarking and analysis. In the second setting, we fix $n_{shot}=2$ and experiment with $n_{task}\in\{8,16,24,32,40,48,56,64\}$. Note that to ensure the in-context examples are balanced and every class is covered, $n_{shot}=2$ refers to using 2 examples per class for in-context learning. In both scaling settings, we are able to effectively create contexts of sizes ranging from 4k to 32k tokens. [^6] It is possible to further increase the context length, e.g., reaching 128k tokens with 64 tasks and 8 shots. Due to compute constraints, we limit the context length to 32k in this work.
We defer additional implementation and engineering details in §A.4.
@@ 5.1 Main Results
Table 2: Main Results: Fixing 16 Tasks, Scaling the Number of Shots. “S-acc” stands for Single-task ICL accuracy averaged over all 16 tasks, and “L-acc” stands for Lifelong ICL accuracy. “pass” represents the “pass rate” defined in §3.2, i.e., percentage of cases that Lifelong ICL is not significantly worse than Single-task ICL among 5 random samples of few-shot training sets. L-acc is expected to be not worse than S-acc, and the pass rate is expected to be close to 100%.
@@
Llama-2-7B (32k) 61.9 69.8 63.3 77.5 72.8 64.5 53.8 75.6 63.0 41.2 78.0 - -
Llama-2-7B (80k) 38.4 47.6 60.0 100.0 49.8 60.2 100.0 56.3 62.3 96.3 59.8 61.5 76.3
Llama-3-8B (1048k) 51.2 65.5 68.1 78.8 70.0 69.1 76.2 71.5 70.1 71.3 73.6 70.1 57.5
Llama-3-70B (1048k) 60.7 79.1 72.9 68.8 79.0 74.4 50.0 80.3 75.3 57.5 81.7 75.7 51.2
Llama-3.1-70B (128k) 58.8 81.7 81.2 80.0 82.8 81.1 76.2 84.6 82.4 83.8 85.2 83.3 80.0
@@
We present the aggregated results (mean accuracy and overall pass rate) of the Scale-Shot setting in Table 2 and the results of the Scale-Task setting in Table 7. The overall pass rates fall below 90% in 50 out of 54 cases reported in Table 2 and in 41 out of 44 cases in Table 7. When scaling to 32k context with 8 shots and 16 tasks, 9 out of the 11 open-weight models achieve pass rates lower than 60%, suggesting that these models are still far from fully utilizing and flexibly conditioning on the provided context. In the most extreme case, Yi-6B (200k) achieves a pass rate of merely 38.8% in the 8-shot (32k) setting.
While model developers commonly use near-perfect needle-in-a-haystack results as evidence of successful long context utilization (01.AI et al., 2024; GradientAI, 2024b, a), our Task Haystack exposes previously unknown limitations of these models and suggest that these models are far from perfect when deeper, contextual understanding is required.
@@
To have a holistic view on this, we visualize the results from our Scale-Shot experiments by plotting the Lifelong ICL accuracy and pass rate as a function of Single-task ICL accuracy in Fig. 3. For nearly all models, pass rates decrease when the context length increases, highlighting that while these models are able to take in long context as inputs, they are not necessarily utilizing them effectively. For model-wise comparison, GPT-4o takes the lead in terms of both the ICL accuracy and the pass rate. Llama-3.1-70B stands out as the leading open-weight model, achieving an average pass rate of 80%, which is close to the 85% pass rates of GPT-4o and Gemini-1.5-Flash.
One outlier that we notice is the Llama-2-7B (80k) model (Fu et al., 2024), which achieves low ICL accuracies but high pass rates. We notice that this model is trained on language modeling objectives without further instruction tuning or RLHF, which may be the reason behind this trend. This observation also suggests that the pass rates should always be considered together with metrics representing the model’s core capabilities.
@@
Task Haystack supports straightforward visualization for diagnosing model vulnerabilities. In Fig. 3 we present the results of Task Haystack (Scale-Shot Setting) in a way similar to the original needle-in-a-haystack (NIAH) evaluation. While FILM-7B achieves near-perfect results in the original NIAH eval, Fig. 3 suggests that it’s vulnerable when the context length exceeds 12k, particularly when the relevant information appears in the first 75% of the context window. We include NIAH-style visualizations for all compared models in Fig. 9-22. In addition, we provide examples of aggregating results by permutations, by depth in the context, and by task in Fig. 25-29. We further discuss our findings in §E.2.